A Complete Bibliography of Publications in Conservation Biology: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.01

Title word cross-reference

1987 [Ano87e]. 1989 [Har89a, Har89b].

Absence [Faj89]. Abundance [PE89]. Acid [SN88]. Acidic [HH89]. Act [SB88a]. Advantages [Bou88]. Advocates [Sal89]. Affected [Old88]. Africa [McC87, RR88]. Agencies [Ore87, Wal88]. Agency [SS87]. Agricultural [BD88]. Agriculture [AAM87, Mcn87]. Agroecodevelopment [Dah87]. Also [Hal87]. Alter [Faj89]. Alteration [Hut89]. Altieri [Bru89a]. Amazonia [Fea87, Ore87, PBKA89]. Ambystoma [HH89]. Ammonia [PBBC87]. Among [LLAB88]. Analysis [DDP88, QWJ89]. Animal [Sco88]. Animals [Sch89]. Announcements [Ano89a]. Annual [Har89a, Har89b]. Any [Goo87b]. Anyone [Dis89, Ehr89a]. Appalachian [SB88b]. Approach [HJW88]. aquatilis [LLAB88]. Aquifer [Inf89]. Archie [Ehr87]. Archipelago [Sch89]. Area [Thi89, VG89]. Arid [Gri89]. Articles [Ano87n]. Asia [Mcn87]. Assessment [CHFG88, HA89, SCEA89]. August [Har89b].

B.P. [OS89]. Bane [RR88]. Bank [Goo87a]. Bats [WLP89]. Beach [HMD^{+}89]. Bear [Hor89]. Berkshire [Hal89]. Between [ZGD88]. Big [Ter88]. Biocultural [Bae89]. Biodiversity [Kau88, Nos89, SCEA89]. Biological [Blo88, Car89a, Des89, Gra88, Hue88, HJW88, Led87, Nor87, Ore87, Pat87, RZ89, SS87]. Biologist [Lov89, Nor88b]. Biologists [Des89, Hal87, Wal88]. Biology [All88, Ano89g, Goo87b, Gra88, Jan89, Sco88, TS89, Ano87f, Ano87i, Ano87j, Ano87k, Ano87l, Ano88f, Ano88g, Ano88h, Ano88i, Ano89i, Ano89j, Ano89k, Ano89f]. Biopsy [DSFT87]. Biosphere [SB88b, Orr89a]. Biotic [Faj89, Har88]. Bird [DBV87, Sal89, TC88, WB89]. Birds [Hut89, RW89, SBA^{+}88b, Thi89]. Bison [ST87]. Black [Cla87, TW88, Hal89]. Black-Footed [Cla87, TW88]. Bond [Old88]. Book [Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89c, Ano89d, Ano89e, Ano89b]. Books [Ano87e]. Boon [RR88]. born [WBNC87]. Bottomland [BCHG89]. Brazilian [Ore87, Fea87]. Breed [CTH89, Hal89]. Breeding [Qui89]. Breeds [Bra88]. Bringing [TS89]. British [CTH89, Hal89]. Bura [Led87]. Butterflies [EM87]. Butterfly [MW88b]. Bylaws [Ano87n].

California [Mil87, Sal89]. Campus [Orr89a]. Can [Mcn87]. Canada [Har89b, WB89]. Captive [ZW87]. Carajás [Ore87]. Carbon [Faj89]. Care [Dis89]. Carr [Ehr87]. Caryophyllaceae [Jen88]. Case [Mil87, PBBC87]. Catalina [RZAM89]. Central [BD88, Sal89]. Change [Gra88]. Changes [BCHG89, Faj89, Yah88]. Changing [Car89a]. Chaparral [SBA^{+}88b]. Chaparral-Requiring [SBA^{+}88b]. Checkerspot [EM87]. Classification [Bou88]. Climatic [Faj89, Gra88]. Climax [Nie87]. Cluster [SB88b]. Coarse [HJW88]. Coarse-Filter [HJW88]. Cockaded [Csu88]. Coexist [Mcn87]. Colonies [dK89]. Comment [Gil88]. Communities [EDR89, WB89, Yah88]. Community [Nie87]. Comparison [Qui89]. Complementary [Wal88]. Concerns [Led87]. Conclusions [Mur89]. Confusion [Mur89]. Consequences [SC87]. Conservation [All88, AL88, AAM87, Alt89, Ano87f, Ano87i, Ano87j, Ano87k, Ano87l, Ano88f, Ano88g, Ano88h, Ano88i, Ano89g, Ano89i, Ano89j, Ano89k, Ano89n, Bru89b, Car89a, Des89, DL89, EEE89, EM87, Gai89, Gei88, Goo87a, Goo87b, Gra88, Gri89, Hal87, Hal89, Har88, Hut89, Lac88, Led87, May88, MV88, Mil87, Mur89, Nor88b, OS89, Pat87, PBKA89, PBBC87, RZ89, Rob89, Sco88, She88, SC87, TB89, Thi89, TS89, TPD89, Wal88, Wil87b, ZGD88, vVH89, Ano89f]. Conserving [Blo88, RW89, Wal88]. Constancy [Nor88a]. Control [Dob88]. Coral [McC87]. Corridors [Nos87, SC87]. Cost [RBT88]. Costs [SC87]. Councilors [Ano87f]. Cox [Nos87]. Cranes [ZW87, ZGD88]. Crisis [Ehr89b]. Critical [TB89]. Critique [LVT88]. Crop [AAM87, Alt89, Bru89b]. Cryptic [PE89]. Cutthroat [AL88]. Cyclic [TB89].

Dams [Mcn87]. Dasyuridae [BL87]. Data [RTS89]. Deborah [Sim88]. Deciduous [Hut89]. Deer [AWS88]. Deforestation [Fea87]. Degradation [McC87]. deltoides [Jen88]. Deposition [HH89]. Desert [EEE89, MV88]. Design [DDP88, Gra88]. Desmognathus [SB88b]. Development [Car87, Dah87, Fea87, Mcn87, She89, SS87]. Dianthus [Jen88]. Differences [ZGD88]. Dioxide [Faj89]. Dipodomys [PE89]. Discharges [Inf89]. Disease [May88, Sco88, TW88]. Dispersal [Sou89, TB89]. Distribution [AL88, PE89]. Disturbed [SB88b]. Diversity [Ano87g, Bae89, Blo88, Dah87, Har88, HJW88, JH88, Lac87, Led87, LLAB88, Nor87, Ore87, RZ89, SS87, She88, Sou87b, SSD88, Wal88]. Do [Goo87b]. Does [Dis89]. Dragon [Kau88]. Drift [Lac87]. Drosophila [FG89]. Drought [TB89]. dwelling [EEE89]. Dynamics [DL89, Nie87, SBA^{+}88b, TC88].

E. [Ter88]. East [McC87]. Ecological [Bou88, LVT88, Orr89b, SB88b]. Ecologists [Hal87]. Ecology [EDR89, Jor88]. Economic [Fea87, PBKA89]. Ecosystem [Hal87, Jor88]. Ecosystems [Bou88, CHFG88, Dob88]. Edge [AWS88, Har88]. Edges [Yah88]. Editorial [Ano87h, Con88, Ehr88, Jan87, Sal87, Wil87a]. Education [Rob89]. Effect [Hut89]. Effective [HA89, ST87]. Effects [AWS88, BD88, Har88, HH89, Jen88, Lac87, Led87, Sch89, SN88]. Endangered [LLAB88, TW88, SB88a]. Endemic [RZAM89, WOG87]. Enriched [Faj89]. Environmental [Ehr89b]. Environments [Car89a, Faj89]. Equivalents [RBT88]. Erratum [Ano88e, Ano89h, SBA^{+}88a]. Estimates [RBT88, WLP89]. Estimating [PE89]. Estimation [HMS87]. Estimators [HA89]. Ethic [Nor88a]. Ethnobotany [PBBC87]. Eurasian [WB89]. Even [Faj89]. Evolutionary [Jan89]. Example [DDP88, TW88]. Examples [Sch89]. Expansion [BD88]. Experimental [QWJ89]. Exploitation [Old88]. Exploration [NM87]. Extant [Dia87, Kin88]. Extinct [Dia87, Kin88]. Extinction [FG89, Gil88, Mye87, QH87, QH88, QWJ89, Ver89a]. Extinctions [EDR89, OS89, SBA^{+}88b, WBNC87]. Extreme [WOG87].

Face [Car89a]. Faunal [BCHG89]. Federal [CHFG88]. Female [Old88]. Female-Offspring [Old88]. Feral [vVH89]. Ferret [Cla87, TW88]. Filter [HJW88]. Financing [Goo87a, Ore87]. Fish [EEE89, RTS89]. Fishes [All88, MV88, She88]. Flies [FG89, RR88]. Florida [TB89]. Footed [Cla87, TW88]. Foraging [ZGD88]. Forest [BCHG89, EDR89, Hut89, Mil87, SN88, Thi89, VG89]. Forest-harvested [VG89]. Forests [AWS88, PBKA89, Whi87]. Fragmentation [Jen88]. Fragmented [TC88]. French [Thi89]. Frequency [BD88]. Fruit [WLP89]. Fruits [VG89]. furbishiae [WOG87].

Galápagos [Sch89]. Gambusia [EEE89]. Game [RW89, Thi89]. General [NM87]. Genetic [AL88, Alt89, Bra88, Bru89b, CTH89, Dah87, Lac87, Mil87, TPD89, WOG87, vVH89]. Genetically [HA89]. Genetics [EEE89, Lac88, MV88, SB88b]. Genic [LLAB88]. Genotypic [RTS89]. Germplasm [Mil87]. Gila [SHP89]. Gill [Sal89]. Gilpin [QH88]. Gloucester [Hal89]. Goals [Ano87i, Ano87j, Ano87k, Ano88f, Ano88g, Ano88h, Ano88i, Ano89i, Ano89j, Ano89k]. Got [Ano87l]. Gradient [DDP88]. Grande [Ore87]. Grizzly [Hor89]. Growth [Doa89]. guadramaculatus [SB88b]. Guiana [Thi89]. Gulf [HMD^{+}89].

Habitat [Hut89, Jen88, QWJ89, SB88a, SBA^{+}88b, TB89, TC88]. Habitat-Interior [TC88]. Habitats [BL87, Gil88, Mcn87, QH87, QH88, SN88]. habroptilus [TPD89]. Hand [Qui89]. Hand- [Qui89]. Hardwood [BCHG89]. Harvest [She89]. harvested [VG89]. Harvesting [Gri89]. Hastings [Gil88]. Hebridean [CTH89]. Here [Ano87l]. Heteromyidae [PE89]. hid [Gri89]. Historical [Ano89g]. History [Ano87l, Ver89a]. Howellia [LLAB88]. Human [Old88]. Hunting [Gei88]. Hybridization [RZAM89].

Immigration [Lac87]. Impact [Sco88]. Impending [Mye87]. Implications [Car89a, Hal89, Pat87, PE89, SN88, Sco88, ZGD88]. Importance [Wil87b]. Important [PBKA89]. Inbreeding [CTH89, RBT88]. Indian [CHFG88]. Indicator [LVT88]. Indonesia [Whi87, WBNC87]. Indonesian [dK89]. Industrial [BD88]. Infection [Sco88]. Influence [TS89]. Interacting [Lac87]. Interior [TC88]. International [Fea87, Hor89, Kin88, Nor87, Ore87, SS87]. Intertidal [QWJ89]. Introduced [Dob88, Sch89, WB89]. Invasion [Ver89a]. Invasions [Car89a, Hue88]. Invertebrates [Wil87b]. Iquitos [VG89]. Irrigation [Led87]. Island [DBV87, Dob88, RZAM89, Sch89]. Islands [HMD^{+}89, SBA^{+}88b, WLP89]. Isolated [DBV87].

Javan [DBV87]. Jeopardize [Gei88].

Kakapo [TPD89]. Kangaroo [Gri89]. Kelp [EDR89]. Kenya [Led87]. Key [HMD^{+}89]. Kite [TB89]. Know [Ver89b].

Lack [LLAB88]. Land [Hor89, Hut89, Nor88a, Wal88]. Land-Use [Hor89]. Lands [CHFG88]. Landscapes [Nos87, TC88]. Large [HA89, Hal89]. Last [Ver89a]. Law [BD88]. Legal [Kin88]. Legislative [Blo88]. Lending [Nor87]. Leopold [Nor88a]. Lessons [EM87, TB89, Goo87b]. Lethal [RBT88]. Letter [Car89b]. Letters [How89, Hum88, Jan88, Len89, Ros87, Sch88]. Levels [She89]. Limitations [Bou88]. Limited [Des89]. Listening [Ehr89a]. Literacy [Orr89b]. Little [Wil87b]. Livestock [vVH89]. Local [Inf89]. Logging [Doa89]. Loghtan [CTH89]. Long [EM87, MW88b]. Long-Term [EM87, MW88b]. Lop [Hal89]. Loss [BCHG89, Lac87, Nor87, Whi87]. Losses [Bra88]. Louisiana [BCHG89].

Mahogany [RZAM89]. Maintaining [Bae89, HJW88]. Major [Bra88, CHFG88, Goo87a, Mcn87]. Mammals [Dob88, HA89, Old88, RBT88]. Man [Car89a]. Managed [Lac87]. Management [MV88, Nie87, PE89, TS89]. Manitoba [WB89]. Manx [CTH89]. Mariana [WLP89]. mariannus [WLP89]. Marine [Ano89n, Kau88, QWJ89]. Markets [Gei88]. Marsh [She89]. Marsupialia [BL87]. May [Faj89]. Means [Goo87a]. Meat [Gei88]. Meeting [Ano87m, Har89b, Har89a]. Megafaunal [OS89]. Membership [Sou87a]. Memoriam [Ehr87, Sim88]. Message [OS89]. Method [DSFT87]. Methods [RTS89]. Mexican [Hut89]. Mice [She89]. Middle [Hal89]. Migratory [Hut89]. Million [Ver89a]. Minimum [HMS87]. Minor [Bra88]. Mirror [DBV87]. Mississippi [ZW87, ZGD88]. Misuse [VG89]. Modeling [TC88]. Monitoring [MW88b]. Mountain [HH89]. Mountains [BD88, ST87]. Mouse [HMD^{+}89]. Mutation [Lac87].

National [HMD^{+}89, Jan89]. Native [WB89, ZGD88]. Natural [BL87, Mcn87, TS89]. Nature [DDP88]. Near [Car89a, Yah88]. Near-Shore [Car89a]. Need [Ver89b]. Nested [Pat87]. Netting [Sal89]. News [Ano88j, Ano88k, Ano88l, Ano88m, Ano89l, Ano89m, Bru87]. Next [Ehr89b]. nobilis [EEE89]. Nondestructive [RTS89]. Norms [NM87]. North [Ver89a]. Northern [AWS88]. Northwest [Doa89]. November [SSD88].

O [Ter88]. Objectives [Ano87i, Ano87j, Ano87k, Ano88f, Ano88g, Ano88h, Ano88i, Ano89j, Ano89k, Ano89i]. Obligations [Lov89]. Observatory [Sal89]. Obtaining [DSFT87, RTS89]. Ocean [Car89a]. Offspring [Old88]. Old [Doa89, Hal89]. Oligarchic [PBKA89]. One [WBNC87]. Ontario [Har89b]. Other [Led87]. Overfishing [McC87]. Owls [Doa89].

Pacific [Doa89]. Paleoecological [Gra88]. Paleoecology [HJW88]. Paraguay [Gai89, RZ89]. Parasites [Dob88]. Parent [ZW87]. Parent-Reared [ZW87]. Parks [Jan89]. Parts [Gei88]. Patch [QWJ89]. Patterns [She88]. Peasant [AAM87]. Pecos [EEE89]. Pedicularis [WOG87]. Pelagic [Car87]. Pelecypod [Ver89a]. Perdido [HMD^{+}89]. Peromyscus [HMD^{+}89]. Persist [Goo87b]. Perspective [Ano89g, Gra88, Hut89]. Perspectives [Car87]. Philosophy [NM87]. Pigs [Hal89]. Plan [MW88b]. Plant [AAM87, LLAB88, Nie87, WB89]. Plants [PBKA89, Sch89]. Plovers [Qui89]. Poeciliidae [EEE89]. Point [Sal89]. Policies [NM87]. Policy [Goo87a]. polionotus [HMD^{+}89]. Pollination [Jen88]. Polymorphisms [CTH89]. Polytypic [AL88]. Population [DL89, FG89, HMS87, HA89, HH89, Hor89, Lac87, Lac88, SST88, ST87, WLP89]. Populations [DL89, Lac87, LLAB88, Sco88, TC88, vVH89]. Portland [CTH89]. Position [TS89]. Possible [HH89]. Potential [Dob88]. Practise [MW88a]. Prairie [WB89]. Precedent [Kin88]. Precipitation [SN88]. Preliminary [CHFG88, McC87, NM87]. Primate [DL89]. Principle [Pat87]. Principles [NM87]. Priorities [Dah87]. Private [Gai89]. Privileges [Gei88]. Program [Hue88, SS87, Whi87]. Programs [Gai89]. Progress [Blo88, Cla87, PP88]. Project [Led87, Mil87]. Projects [Fea87, Mcn87]. Protection [Bou88, SCEA89, SB88a]. Proven [Dia87, Kin88]. Psittaciformes [TPD89]. Pteropus [WLP89]. Public [Wal88].

Quantitative [PBBC87]. Quinn [Gil88].

Rabinowitz [Sim88]. Rain [Thi89, Whi87]. Rangelands [Gri89]. Rapid [SBA^{+}88b]. Raptors [RW89, Thi89]. Rare [CTH89, Hal89]. Rarity [BL87, She88]. Real [Nos87]. Reared [Qui89, ZW87]. Received [Ano87e, Ano89b]. Recharge [Inf89]. Reconstructed [SBA^{+}88b]. Recovery [Cla87, PP88, SHP89]. Recreational [BD88]. Recruitment [Sou89]. Red [Csu88, PP88]. Red-Cockaded [Csu88]. Redefining [Dah87]. Reef [McC87]. Reestablishment [HMD^{+}89]. Refuge [ZW87, ST87]. Regeneration [Orr89c]. Rejoinder [Bru89a]. Relation [BL87, Nie87]. Relationships [CTH89]. Released [ZW87, ZGD88]. Remote [DSFT87]. Reply [Nos87, QH88]. Report [Ano87m, Cla87, Hue88, Lac88, McC87, PP88]. Requirements [Thi89]. Requiring [SBA^{+}88b]. Research [MW88a]. Reserve [DDP88, SB88b]. Reserves [Gra88]. Resource [Alt89, Bru89b, Mcn87, PBKA89, TS89]. Resources [AAM87]. Responsibility [Wal88]. Restoring [Dob88]. Rethinking [Alt89, Bru89b]. Reviews [Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89c, Ano89d, Ano89e]. Reyes [Sal89]. Rising [She89]. Risk [She88]. Rocky [BD88, HH89]. Rodentia [PE89]. Role [Car89a, Des89, Gra88, Whi87]. Roles [Wal88]. Run [Ter88, Wil87b].

Saddleback [Hal89]. Salamander [HH89, SB88b]. Sale [Gei88]. Salt [She89]. Sample [HMS87]. Samples [DSFT87]. Sandhill [ZW87, ZGD88]. Saving [Ver89b]. Scale [Mur89]. Scientists [Sal89]. SCOPE [Hue88]. Scrophulariaceae [WOG87]. Sea [Car87, She89, Ver89b, Ver89a]. Seabird [dK89]. Seashore [HMD^{+}89]. Section [Ano89n]. Seed [Jen88]. Selected [NM87]. Selection [Lac87]. Semi [Gri89]. Semi-hid [Gri89]. September [Ano87e]. Sequel [Ter88]. Set [Jen88]. Settlement [Led87]. Sharing [Wal88]. Sheep [CTH89]. Shore [Car89a]. Simberloff [Nos87]. Simulation [SST88]. Size [HA89, QWJ89, ST87]. Sizes [HMS87]. Skin [DSFT87]. Sleeping [Kau88]. Sminthopsis [BL87]. Snail [QWJ89, TB89]. Snowy [Qui89]. Society [Ano87f, Ano87i, Ano87j, Ano87k, Ano87l, Ano87n, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Bru87]. South [Alt89]. Southeastern [Ore87]. Southern [SB88b]. Spasm [Mye87]. Spatial [FG89]. Speak [Nos89]. Special [Ano89n]. Species [AL88, Goo87b, LVT88, Mil87, Mur89, PE89, SB88a, TW88]. Spot [Hal89].

Spotted [Doa89]. Spring [EEE89, Inf89]. Spring-dwelling [EEE89]. Stabilize [Inf89]. Stage [Car87]. States [CHFG88]. Status [CHFG88, Hue88, SCEA89]. stephensi [PE89]. Stochastic [SST88]. Story [SHP89]. Stream [She88]. Strigops [TPD89]. Structure [CTH89, FG89]. Structures [Hal89]. Studies [EM87]. Study [FG89]. Subdivided [Gil88, QH87, QH88]. Subdivision [Lac87, QWJ89]. Subsets [Pat87]. Success [SHP89]. Succession [Nie87]. Sulawesi [WBNC87]. Survival [DBV87, ZW87]. Synergisms [Mye87]. Systems [Faj89].

Tamworth [Hal89]. Task [Orr89d]. Temperate [Mil87]. Tensas [BCHG89]. Term [EM87, MW88b]. Terrestrial [CHFG88]. Their [CTH89]. Things [Ter88, Wil87b]. Threatened [MW88b, Old88]. Threatens [Hor89]. Threats [dK89]. Three [Ver89a]. Tiger [HH89]. tigrinum [HH89]. Too [AWS88]. Topminnow [SHP89]. Toronto [Har89b]. Transmigration [Whi87]. Tree [Mil87]. trissyllepsis [HMD^{+}89]. Tropical [Hut89, Mcn87, PBKA89, RW89, Whi87]. Trout [AL88]. Tsetse [RR88]. Turtle [Car87].

U.S. [Blo88, SS87]. Uncontrolled [Hor89]. United [CHFG88]. University [Orr89c, Har89b]. Unless [Dia87, Kin88]. Unnatural [BL87]. Urban [She89, SBA^{+}88b]. USA [BD88]. Use [Hor89, VG89]. Uses [LVT88]. Utilization [PBKA89].

Variation [AL88, TPD89, WOG87]. Vegetation [Nie87, Sch89, WB89]. Vertebrate [LVT88]. Viable [HMS87]. View [Alt89]. Violations [BD88]. virginiae [BL87]. Visitation [Jen88]. vs [MW88a].

Water [Mcn87]. Watershed [BCHG89]. Watersheds [SB88b]. West [Hut89]. Wetland [CHFG88]. White [Hal89]. Who [Nos89]. Wichita [ST87]. Wild [AAM87, Qui89]. Wild-Reared [Qui89]. Wildlands [Goo87a]. Wildlife [BD88, Gei88, Mcn87, SN88, ST87, Yah88]. Will [Nos89]. Wilson [Ter88]. Wisconsin [AWS88]. Within [LLAB88]. Wolf [NM87, PP88]. Woodland [DBV87]. Woodpecker [Csu88]. Work [Mye87]. World [Ter88, Wil87b, Goo87a]. Wrong [Mur89].

Years [OS89, Ver89a].

Zealand [DDP88]. Zoo [Rob89].

References

Altieri:1987:PAC

[AAM87] Miguel A. Altieri, M. Kat Anderson, and Laura C. Merrick. Peasant agriculture and the conservation of crop and wild plant resources. Conservation Biology, 1(1):49–58, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:1988:CDG

[AL88] Fred W. Allendorf and Robb F. Leary. Conservation and distribution of genetic variation in a polytypic species, the cutthroat trout. Conservation Biology, 2(2):170–184, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:1988:CBF

[All88] Fred W. Allendorf. Conservation biology of fishes. Conservation Biology, 2(2):145–148, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Altieri:1989:RCG

[Alt89] Miguel A. Altieri. Rethinking crop genetic resource conservation: a view from the South. Conservation Biology, 3(1):77–79, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See rejoinder [Bru89a].

Anonymous:1987:BRa

[Ano87a] Anonymous. Book reviews. Conservation Biology, 1(1):82–85, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:BRb

[Ano87b] Anonymous. Book reviews. Conservation Biology, 1(2):175–180, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:BRc

[Ano87c] Anonymous. Book reviews. Conservation Biology, 1(3):265–269, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:BRd

[Ano87d] Anonymous. Book reviews. Conservation Biology, 1(4):347–352, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:BRT

[Ano87e] Anonymous. Books received through September 1987. Conservation Biology, 1(4):352–353, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:CSC

[Ano87f] Anonymous. Councilors of the Society for Conservation Biology. Conservation Biology, 1(2):183, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:D

[Ano87g] Anonymous. Diversity. Conservation Biology, 1(1):80–86, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:E

[Ano87h] Anonymous. Editorial. Conservation Biology, 1(1):6–7, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:GOSa

[Ano87i] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 1(1):91, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:GOSb

[Ano87j] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 1(3):270, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:GOSc

[Ano87k] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 1(4):354, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:HSC

[Ano87l] Anonymous. History of the Society for Conservation Biology: How and why we got here. Conservation Biology, 1(1):4–5, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:MR

[Ano87m] Anonymous. Meeting report. Conservation Biology, 1(3):186–187, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1987:SAB

[Ano87n] Anonymous. Society articles and bylaws. Conservation Biology, 1 (1):87–90, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:BRa

[Ano88a] Anonymous. Book reviews. Conservation Biology, 2(1):124–125, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:BRb

[Ano88b] Anonymous. Book reviews. Conservation Biology, 2(2):222–228, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:BRc

[Ano88c] Anonymous. Book reviews. Conservation Biology, 2(3):300–301, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:BRd

[Ano88d] Anonymous. Book reviews. Conservation Biology, 2(4):404–409, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:E

[Ano88e] Anonymous. Erratum. Conservation Biology, 2(3):244, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:GOSa

[Ano88f] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 2(1):127, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:GOSb

[Ano88g] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 2(2):229, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:GOSc

[Ano88h] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 2(3):304, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:GOSd

[Ano88i] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 2(4):410, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:NSa

[Ano88j] Anonymous. News of the Society. Conservation Biology, 2(1):3, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:NSb

[Ano88k] Anonymous. News of the Society. Conservation Biology, 2(2):131, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:NSc

[Ano88l] Anonymous. News of the Society. Conservation Biology, 2(3):235– 236, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1988:NSd

[Ano88m] Anonymous. News of the Society. Conservation Biology, 2(4):305– 306, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:A

[Ano89a] Anonymous. Announcements. Conservation Biology, 3(1):16–18, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:BRd

[Ano89b] Anonymous. Book received. Conservation Biology, 3(4):421–423, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:BRa

[Ano89c] Anonymous. Book reviews. Conservation Biology, 3(2):204–210, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:BRb

[Ano89d] Anonymous. Book reviews. Conservation Biology, 3(3):322–327, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:BRc

[Ano89e] Anonymous. Book reviews. Conservation Biology, 3(4):416–423, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:CB

[Ano89f] Anonymous. Conservation Biology. Conservation Biology, 3(1):6, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:CBH

[Ano89g] Anonymous. Conservation biology in historical perspective. Conservation Biology, 3(1):99–106, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:E

[Ano89h] Anonymous. Erratum. Conservation Biology, 3(1):9, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:GOSa

[Ano89i] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 3(2):211, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:GOSb

[Ano89j] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 3(3):328, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:GOSc

[Ano89k] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 3(4):424, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:NSa

[Ano89l] Anonymous. News of the Society. Conservation Biology, 3(1):4, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:NSb

[Ano89m] Anonymous. News of the Society. Conservation Biology, 3(2):113– 115, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1989:SSM

[Ano89n] Anonymous. Special section: Marine conservation. Conservation Biology, 3(3):239, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alverson:1988:FTD

[AWS88] William S. Alverson, Donald M. Waller, and Stephen L. Solheim. Forests too deer: Edge effects in northern Wisconsin. Conservation Biology, 2(4):348–358, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baer:1989:MBD

[Bae89] Adela Baer. Maintaining biocultural diversity. Conservation Biology, 3(1):97–98, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burdick:1989:FCB

[BCHG89] David M. Burdick, Douglas Cushman, Robert Hamilton, and James G. Gosselink. Faunal changes and bottomland hardwood forest loss in the Tensas Watershed, Louisiana. Conservation Biology, 3(3):282–292, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1988:EAI

[BD88] Joel Berger and Dennis Daneke. Effects of agricultural, industrial, and recreational expansion on frequency of wildlife law violations in the Central Rocky Mountains, USA. Conservation Biology, 2 (3):283–289, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Braithwaite:1987:RSV

[BL87] R. W. Braithwaite and W. M. Lonsdale. The rarity of Sminthopsis virginiae, (Marsupialia: Dasyuridae) in relation to natural and unnatural habitats. Conservation Biology, 1(4):341–343, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blockstein:1988:ULP

[Blo88] David E. Blockstein. U.S. legislative progress toward conserving biological diversity. Conservation Biology, 2(4):311–315, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bourgeron:1988:ALE

[Bou88] Patrick S. Bourgeron. Advantages and limitations of ecological classification for the protection of ecosystems. Conservation Biology, 2 (2):218–220, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bratton:1988:MBM

[Bra88] Susan P. Bratton. Minor breeds and major genetic losses. Conservation Biology, 2(3):297–299, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:1987:NS

[Bru87] Peter F. Brussard. News of the Society. Conservation Biology, 1(4): 274, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brush:1989:RA

[Bru89a] Stephen B. Brush. Rejoinder to Altieri. Conservation Biology, 3 (1):80–81, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Alt89].

Brush:1989:RCG

[Bru89b] Stephen B. Brush. Rethinking crop genetic resource conservation. Conservation Biology, 3(1):19–29, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carr:1987:NPP

[Car87] Archie Carr. New perspectives on the pelagic stage of sea turtle development. Conservation Biology, 1(2):103–121, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carlton:1989:MRC

[Car89a] James T. Carlton. Man’s role in changing the face of the ocean: Biological invasions and implications for conservation of near-shore environments. Conservation Biology, 3(3):265–273, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carr:1989:L

[Car89b] Archie Carr III. Letter. Conservation Biology, 3(4):332–333, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crumpacker:1988:PAS

[CHFG88] David W. Crumpacker, Stephen W. Hodge, Dale Friedley, and William P. Gregg, Jr. A preliminary assessment of the status of major terrestrial and wetland ecosystems on Federal and Indian lands in the United States. Conservation Biology, 2(1):103–115, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:1987:BFF

[Cla87] Tim W. Clark. Black-footed ferret recovery: a progress report. Conservation Biology, 1(1):8–13, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Conway:1988:E

[Con88] William Conway. Editorial. Conservation Biology, 2(2):132–134, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Csuti:1988:RCW

[Csu88] Blair Csuti. Red-cockaded woodpecker. Conservation Biology, 2(2): 136–137, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clarke:1989:GPT

[CTH89] S. W. Clarke, E. M. Tucker, and S. J. G. Hall. Genetic polymorphisms and their relationships with inbreeding and breed structure in rare British sheep: The Portland, Manx Loghtan, and Hebridean. Conservation Biology, 3(4):381–388, December 1989. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dahlberg:1987:RDP

[Dah87] Kenneth A. Dahlberg. Redefining development priorities: Genetic diversity and agroecodevelopment. Conservation Biology, 1(4):311– 322, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diamond:1987:BSI

[DBV87] Jared M. Diamond, K. David Bishop, and S. Van Balen. Bird survival in an isolated Javan woodland: Island or mirror? Conservation Biology, 1(2):132–142, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Develice:1988:GAN

[DDP88] Robert L. Develice, Jean Ward Develice, and Geoff N. Park. Gradient analysis in nature reserve design: a New Zealand example. Conservation Biology, 2(2):206–217, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deshmukh:1989:LRB

[Des89] Ian Deshmukh. On the limited role of biologists in biological conservation. Conservation Biology, 3(3):321, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diamond:1987:EUP

[Dia87] Jared M. Diamond. Extant unless proven extinct? Or, extinct unless proven extant? Conservation Biology, 1(1):77–79, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Disney:1989:DAC

[Dis89] R. H. L. Disney. Does anyone care? Conservation Biology, 3(4): 414, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

deKorte:1989:TIS

[dK89] J. de Korte. Threats to Indonesian seabird colonies. Conservation Biology, 3(4):336–340, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobson:1989:PDC

[DL89] Andrew P. Dobson and Anna Marie Lees. The population dynamics and conservation of primate populations. Conservation Biology, 3 (4):362–380, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doak:1989:SOO

[Doa89] Daniel Doak. Spotted owls and old growth logging in the Pacific Northwest. Conservation Biology, 3(4):389–396, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobson:1988:RIE

[Dob88] A. P. Dobson. Restoring island ecosystems: The potential of parasites to control introduced mammals. Conservation Biology, 2(1): 31–39, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dvm:1987:RMO

[DSFT87] Willlam B. Karesh Dvm, Fellner Smith, and Harmony Frazier-Taylor. A remote method for obtaining skin biopsy samples. Conservation Biology, 1(3):261–262, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estes:1989:EEK

[EDR89] James A. Estes, David O. Duggins, and Galen B. Rathbun. The ecology of extinctions in kelp forest communities. Conservation Biology, 3(3):252–264, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Echelle:1989:CGS

[EEE89] Alice F. Echelle, Anthony A. Echelle, and David R. Edds. Conservation genetics of a spring-dwelling desert fish, the Pecos Gambusia (Gambusia nobilis, Poeciliidae). Conservation Biology, 3(2): 159–169, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1987:ACM

[Ehr87] David Ehrenfeld. Archie Carr: In memoriam. Conservation Biology, 1(2):169–172, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1988:E

[Ehr88] David Ehrenfeld. Editorial. Conservation Biology, 2(1):4–5, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1989:AL

[Ehr89a] David Ehrenfeld. Is anyone listening? Conservation Biology, 3(4): 415, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1989:NEC

[Ehr89b] David Ehrenfeld. The next environmental crisis. Conservation Biology, 3(1):1–3, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrlich:1987:CLL

[EM87] Paul R. Ehrlich and Dennis D. Murphy. Conservation lessons from long-term studies of checkerspot butterflies. Conservation Biology, 1(2):122–131, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fajer:1989:HEC

[Faj89] Eric D. Fajer. How enriched carbon dioxide environments may alter biotic systems even in the absence of climatic changes. Conservation Biology, 3(3):318–320, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fearnside:1987:DIE

[Fea87] Philip M. Fearnside. Deforestation and international economic development projects in Brazilian Amazonia. Conservation Biology, 1(3):214–221, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forney:1989:SSP

[FG89] Karin A. Forney and Michael E. Gilpin. Spatial structure and population extinction: a study with Drosophila flies. Conservation Biology, 3(1):45–51, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaito:1989:PCP

[Gai89] Raúl Gaito. Private conservation programs in Paraguay. Conservation Biology, 3(2):120–122, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geist:1988:HMW

[Gei88] Valerius Geist. How markets in wildlife meat and parts, and the sale of hunting privileges, jeopardize wildlife conservation. Conservation Biology, 2(1):15–26, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilpin:1988:CQH

[Gil88] Michael E. Gilpin. A comment on Quinn and Hastings: Extinction in subdivided habitats. Conservation Biology, 2(3):290–292, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [QH87, QH88].

Goodland:1987:WBW

[Goo87a] R. J. A. Goodland. The World Bank’s wildlands policy: a major new means of financing conservation. Conservation Biology, 1(3): 210–213, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodman:1987:HDS

[Goo87b] Daniel Goodman. How do any species persist? Lessons for conservation biology. Conservation Biology, 1(1):59–62, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Graham:1988:RCC

[Gra88] Russell W. Graham. The role of climatic change in the design of biological reserves: The paleoecological perspective for conservation biology. Conservation Biology, 2(4):391–394, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grigg:1989:KHC

[Gri89] Gordon Grigg. Kangaroo harvesting and the conservation of arid and semi-hid rangelands. Conservation Biology, 3(2):194–197, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:1989:GEP

[HA89] Richard B. Harris and Fred W. Allendorf. Genetically effective population size of large mammals: an assessment of estimators. Conservation Biology, 3(2):181–191, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:1987:EEC

[Hal87] Charles A. S. Hall. Ecosystem ecologists are also conservation biologists. Conservation Biology, 1(3):263–264, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:1989:BSC

[Hal89] S. J. G. Hall. Breed structures of for conservation of middle rare pigs: Implications the Berkshire, Tamworth, White, Large Black, Gloucester Old Spot, British Saddleback, and British Lop. Conservation Biology, 3(1):30–38, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:1988:EEC

[Har88] Larry D. Harris. Edge effects and conservation of biotic diversity. Conservation Biology, 2(4):330–332, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:1989:AM

[Har89a] Larry D. Harris. The 1989 Annual Meeting. Conservation Biology, 3(4):331, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:1989:AMU

[Har89b] Larry D. Harris. 1989 annual meeting: University of Toronto, Toronto, Ontario, Canada, August 6–10, 1989. Conservation Biology, 3(1):5–6, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harte:1989:PEA

[HH89] John Harte and Erika Hoffman. Possible effects of acidic deposition on a Rocky Mountain population of the tiger salamander Ambystoma tigrinum. Conservation Biology, 3(2):149–158, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1988:PCF

[HJW88] Malcolm L. Hunter, Jr., George L. Jacobson, Jr., and Thompson Webb III. Paleoecology and the coarse-filter approach to maintaining biological diversity. Conservation Biology, 2(4):375–385, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holier:1989:RPK

[HMD^{+}89] N. R. Holier, D. W. Mason, R. M. Dawson, T. Simons, and M. C. Wooten. Reestablishment of the Perdido Key beach mouse (Peromyscus polionotus trissyllepsis) on Gulf Islands National Seashore. Conservation Biology, 3(4):397–404, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:1987:SSM

[HMS87] Richard B. Harris, Lynn A. Maguire, and Mark L. Shaffer. Sample sizes for minimum viable population estimation. Conservation Biology, 1(1):72–76, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Horejsi:1989:ULU

[Hor89] Brian L. Horejsi. Uncontrolled land-use threatens an international grizzly bear population. Conservation Biology, 3(3):220–226, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Howard:1989:L

[How89] James H. Howard. Letters. Conservation Biology, 3(1):7, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huenneke:1988:SPB

[Hue88] Laura Huenneke. SCOPE program on biological invasions: a status report. Conservation Biology, 2(1):8–14, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Humphrey:1988:L

[Hum88] Stephen R. Humphrey. Letters. Conservation Biology, 2(3):240– 244, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutto:1989:EHA

[Hut89] Richard L. Hutto. The effect of habitat alteration on migratory land birds in a West Mexican tropical deciduous forest: a conservation perspective. Conservation Biology, 3(2):138–148, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Inferrera:1989:LAR

[Inf89] Jeffrey J. Inferrera. Local aquifer recharge to stabilize spring discharges. Conservation Biology, 3(4):413, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Janzen:1987:E

[Jan87] Daniel H. Janzen. Editorial. Conservation Biology, 1(2):95–96, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Janzen:1988:L

[Jan88] Daniel H. Janzen. Letters. Conservation Biology, 2(4):309–310, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Janzen:1989:EBN

[Jan89] Daniel H. Janzen. The evolutionary biology of national parks. Conservation Biology, 3(2):109–112, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jennersten:1988:PDD

[Jen88] Ola Jennersten. Pollination in Dianthus deltoides (Caryophyllaceae): Effects of habitat fragmentation on visitation and seed set. Conservation Biology, 2(4):359–366, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:1988:D

[JH88] Susan K. Jacobson and Jeffrey L. Hardesty. Diversity. Conservation Biology, 2(2):221, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jordan:1988:EE

[Jor88] Carl F. Jordan. Ecosystem ecology. Conservation Biology, 2(2): 137–138, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1988:MBS

[Kau88] Les Kaufman. Marine biodiversity: The sleeping dragon. Conservation Biology, 2(4):307–308, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

King:1988:EUP

[Kin88] F. Wayne King. Extant unless proven extinct: The international legal precedent. Conservation Biology, 2(4):395–397, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lacy:1987:LGD

[Lac87] Robert C. Lacy. Loss of genetic diversity from managed populations: Interacting effects of drift, mutation, immigration, selection, and population subdivision. Conservation Biology, 1(2):143–158, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lacy:1988:RPG

[Lac88] Robert C. Lacy. A report on population genetics in conservation. Conservation Biology, 2(3):245–248, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ledec:1987:EKB

[Led87] George Ledec. Effects of Kenya’s Bura Irrigation Settlement Project on biological diversity and other conservation concerns. Conservation Biology, 1(3):247–258, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lennartz:1989:L

[Len89] Michael R. Lennartz. Letters. Conservation Biology, 3(1):7–9, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lesica:1988:LGD

[LLAB88] Peter Lesica, Robb F. Leary, Fred W. Allendorf, and David E. Bilderback. Lack of genic diversity within and among populations of an endangered plant, Howellia aquatilis. Conservation Biology, 2(3):275–282, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:1989:OB

[Lov89] Thomas Lovejoy. The obligations of a biologist. Conservation Biology, 3(4):329–330, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Landres:1988:EUV

[LVT88] Peter B. Landres, Jared Verner, and Jack Ward Thomas. Ecological uses of vertebrate indicator species: a critique. Conservation Biology, 2(4):316–328, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

May:1988:CD

[May88] Robert M. May. Conservation and disease. Conservation Biology, 2 (1):28–30, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClenahan:1987:OCR

[McC87] Timothy R. McClenahan. Overfishing and coral reef degradation: a preliminary report from East Africa. Conservation Biology, 1(2): 97–102, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcneely:1987:HDW

[Mcn87] Jeffrey A. Mcneely. How dams and wildlife can coexist: Natural habitats, agriculture, and major water resource development projects in tropical Asia. Conservation Biology, 1(3):228–238, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Millar:1987:CFG

[Mil87] Constance I. Millar. The California Forest Germplasm Conservation Project: a case for genetic conservation of temperate tree species. Conservation Biology, 1(3):191–197, October 1987. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:1989:CCW

[Mur89] Dennis D. Murphy. Conservation and confusion: Wrong species, wrong scale, wrong conclusions. Conservation Biology, 3(1):82–84, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1988:CGM

[MV88] Gary K. Meffe and Robert C. Vrijenhoek. Conservation genetics in the management of desert fishes. Conservation Biology, 2(2): 157–169, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:1988:RVP

[MW88a] Ronald I. Miller and Rodger Waldman. Research vs practise. Conservation Biology, 2(2):135, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:1988:LTM

[MW88b] Dennis D. Murphy and Stuart B. Weiss. A long-term monitoring plan for a threatened butterfly. Conservation Biology, 2(4):367– 374, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Myers:1987:ESI

[Mye87] Norman Myers. The extinction spasm impending: Synergisms at work. Conservation Biology, 1(1):14–21, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niering:1987:VDS

[Nie87] William A. Niering. Vegetation dynamics (succession and climax) in relation to plant community management. Conservation Biology, 1(4):287–295, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naess:1987:PWP

[NM87] Arne Naess and Ivar Mysterud. Philosophy of wolf policies I: General principles and preliminary exploration of selected norms. Conservation Biology, 1(1):22–34, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norse:1987:ILL

[Nor87] Elliott A. Norse. International lending and the loss of biological diversity. Conservation Biology, 1(3):259–260, October 1987. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:1988:CLL

[Nor88a] Bryan G. Norton. The constancy of Leopold’s land ethic. Conservation Biology, 2(1):93–102, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:1988:WCB

[Nor88b] Bryan G. Norton. What is a conservation biologist? Conservation Biology, 2(3):237–238, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1987:CRL

[Nos87] Reed F. Noss. Corridors in real landscapes: a reply to Simberloff and Cox. Conservation Biology, 1(2):159–164, August 1987. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SC87].

Noss:1989:WWS

[Nos89] Reed F. Noss. Who will speak for biodiversity? Conservation Biology, 3(2):202–203, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oldfield:1988:TMA

[Old88] Margery L. Oldfield. Threatened mammals affected by human exploitation of the female-offspring bond. Conservation Biology, 2 (3):260–274, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oren:1987:GCI

[Ore87] David C. Oren. Grande Carajás, international financing agencies, and biological diversity in Southeastern Brazilian Amazonia. Conservation Biology, 1(3):222–227, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1989:CB

[Orr89a] David W. Orr. The campus and the Biosphere. Conservation Biology, 3(2):116–117, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1989:EL

[Orr89b] David W. Orr. Ecological literacy. Conservation Biology, 3(4):334– 335, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1989:RU

[Orr89c] David W. Orr. Regeneration of the university. Conservation Biology, 3(3):218–219, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1989:T

[Orr89d] David W. Orr. The task. Conservation Biology, 3(1):10, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Owen-Smith:1989:MEC

[OS89] Norman Owen-Smith. Megafaunal extinctions: The conservation message from 11,000 years B.P. Conservation Biology, 3(4):405– 412, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Patterson:1987:PNS

[Pat87] Bruce D. Patterson. The principle of nested subsets and its implications for biological conservation. Conservation Biology, 1(4):323– 334, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prance:1987:QEC

[PBBC87] G. T. Prance, W. Balée, B. M. Boom, and R. L. Carneiro. Quantitative ethnobotany and the case for conservation in ammonia. Conservation Biology, 1(4):296–310, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peters:1989:OFE

[PBKA89] Charles M. Peters, Michael J. Balick, Francis Kahn, and Anthony B. Anderson. Oligarchic forests of economic plants in Amazonia: Utilization and conservation of an important tropical resource. Conservation Biology, 3(4):341–349, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:1989:EDA

[PE89] M. V. Price and P. R. Endo. Estimating the distribution and abundance of a cryptic species, Dipodomys stephensi (Rodentia: Heteromyidae), and implications for management. Conservation Biology, 3(3):293–301, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:1988:RWR

[PP88] Michael K. Phillips and Warren T. Parker. Red wolf recovery: a progress report. Conservation Biology, 2(2):139–144, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quinn:1987:ESH

[QH87] James F. Quinn and Alan Hastings. Extinction in subdivided habitats. Conservation Biology, 1(3):198–209, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comment [Gil88] and reply [Gil88].

Quinn:1988:ESH

[QH88] James F. Quinn and Alan Hastings. Extinction in subdivided habitats: Reply to Gilpin. Conservation Biology, 2(3):293–296, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gil88].

Quinn:1989:CBH

[Qui89] Patricia L. Quinn. Comparison of the breeding of hand- and wildreared snowy plovers. Conservation Biology, 3(2):198–201, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quinn:1989:EAP

[QWJ89] James F. Quinn, Carole L. Wolin, and Michael L. Judge. An experimental analysis of patch size, habitat subdivision, and extinction in a marine intertidal snail. Conservation Biology, 3(3):242–251, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ralls:1988:ELE

[RBT88] Katherine Ralls, Jonathan D. Ballou, and Alan Templeton. Estimates of lethal equivalents and the cost of inbreeding in mammals. Conservation Biology, 2(2):185–193, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:1989:ZEC

[Rob89] Michael H. Robinson. The zoo that is not: Education for conservation. Conservation Biology, 3(3):213–215, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenblum:1987:L

[Ros87] Edwin E. Rosenblum. Letters. Conservation Biology, 1(4):278–279, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rogers:1988:TFA

[RR88] David J. Rogers and Sarah E. Randolph. Tsetse flies in Africa: Bane or boon? Conservation Biology, 2(1):57–65, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robbins:1989:NMO

[RTS89] Lynn W. Robbins, Deana K. Tolliver, and Michael H. Smith. Nondestructive methods for obtaining genotypic data from fish. Conservation Biology, 3(1):88–91, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:1989:CTR

[RW89] Scott K. Robinson and David S. Wilcove. Conserving tropical raptors and game birds. Conservation Biology, 3(2):192–193, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rios:1989:CBD

[RZ89] Eustacio Ríos and Elsa Zardini. Conservation of biological diversity in Paraguay. Conservation Biology, 3(2):118–120, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rieseberg:1989:HIE

[RZAM89] Loren H. Rieseberg, Scott Zona, Linda Aberbom, and Terrence D. Martin. Hybridization in the island endemic, Catalina mahogany. Conservation Biology, 3(1):52–58, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salwasser:1987:E

[Sal87] Hal Salwasser. Editorial. Conservation Biology, 1(4):275–277, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salzman:1989:SAP

[Sal89] Jamese. Salzman. Scientists as advocates: The Point Reyes Bird Observatory and gill netting in Central California. Conservation Biology, 3(2):170–180, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sidle:1988:HPU

[SB88a] John G. Sidle and David B. Bowman. Habitat protection under the Endangered Species Act. Conservation Biology, 2(1):116–118, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stiven:1988:EGS

[SB88b] Alan E. Stiven and Richard C. Bruce. Ecological genetics of the salamander Desmognathus guadramaculatus from disturbed watersheds in the Southern Appalachian Biosphere Reserve Cluster. Conservation Biology, 2(2):194–205, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1988:E

[SBA^{+}88a] Michael E. Soulé, Douglas T. Bolger, Allison C. Alberts, Raymond Sauvajot, John Wright, Marina Sorice, and Scott Hill. Erratum. Conservation Biology, 2(2):134, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1988:RDR

[SBA^{+}88b] Michael E. Soulé, Douglas T. Bolger, Allison C. Alberts, John Wrights, Marina Sorice, and Scott Hill. Reconstructed dynamics of rapid extinctions of chaparral-requiring birds in urban habitat islands. Conservation Biology, 2(1):75–92, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simberloff:1987:CCC

[SC87] Daniel Simberloff and James Cox. Consequences and costs of conservation corridors. Conservation Biology, 1(1):63–71, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Nos87].

Scott:1989:SAB

[SCEA89] J. Michael Scott, Blair Csuti, J. E. Estes, and H. Anderson. Status assessment of biodiversity protection. Conservation Biology, 3(1):

85–87, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schaff:1988:L

[Sch88] C. Dietrich Schaff. Letters. Conservation Biology, 2(3):239–240, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schofield:1989:EIP

[Sch89] Eileen K. Schofield. Effects of introduced plants and animals on island vegetation: Examples from Galápagos Archipelago. Conservation Biology, 3(3):227–239, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:1988:IID

[Sco88] Marilyn E. Scott. The impact of infection and disease on animal populations: Implications for conservation biology. Conservation Biology, 2(1):40–56, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sheldon:1988:CSF

[She88] Andrew L. Sheldon. Conservation of stream fishes: Patterns of diversity, rarity, and risk. Conservation Biology, 2(2):149–156, June 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shellhammer:1989:SMH

[She89] Howard S. Shellhammer. Salt marsh harvest mice, urban development, and rising sea levels. Conservation Biology, 3(1):59–65, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simons:1989:RGT

[SHP89] Lee H. Simons, Dean A. Hendrickson, and Diana Papoulias. Recovery of the Gila topminnow: a success story? Conservation Biology, 3(1):11–15, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simberloff:1988:DRM

[Sim88] Daniel Simberloff. Deborah Rabinowitz: In memoriam. Conservation Biology, 2(1):119–120, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schreiber:1988:APE

[SN88] R. Kent Schreiber and James R. Newman. Acid precipitation effects on forest habitats: Implications for wildlife. Conservation Biology, 2(3):249–259, September 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1987:M

[Sou87a] Michael E. Soulé. Membership. Conservation Biology, 1(2):94, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1987:D

[Sou87b] Michael Ellman Soulé. Diversity. Conservation Biology, 1(2):173– 174, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1989:DR

[Sou89] Michael Ellman Soulé. Dispersal and recruitment. Conservation Biology, 3(3):216–217, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:1987:BDP

[SS87] Mark L. Shaffer and Kathryn A. Saterson. The biological diversity program of the U.S. Agency for International Development. Conservation Biology, 1(4):280–286, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1988:DN

[SSD88] Michael Soulé, Michael E. Soulé, and Jared Diamond. Diversity November 12, 1987. Conservation Biology, 2(1):121–123, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:1988:SPS

[SST88] Mark L. Shaffer, Alisa M. Shull, and Alan R. Tipton. Stochastic population simulation. Conservation Biology, 2(1):6–7, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shull:1987:EPS

[ST87] Alisa M. Shull and Alan R. Tipton. Effective population size of bison on the Wichita Mountains Wildlife Refuge. Conservation Biology, 1(1):35–41, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Takekawa:1989:CDD

[TB89] Jean E. Takekawa and Steven R. Beissinger. Cyclic drought, dispersal, and the conservation of the snail kite in Florida: Lessons in critical habitat. Conservation Biology, 3(3):302–311, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1988:MDH

[TC88] Stanley A. Temple and John R. Cary. Modeling dynamics of habitat-interior bird populations in fragmented landscapes. Conservation Biology, 2(4):340–347, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Terborgh:1988:BTR

[Ter88] John Terborgh. The big things that run the world — a sequel to E. O. Wilson. Conservation Biology, 2(4):402–403, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thiollay:1989:ARC

[Thi89] J. M. Thiollay. Area requirements for the conservation of rain forest raptors and game birds in French Guiana. Conservation Biology, 3 (2):128–137, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Triggs:1989:GVC

[TPD89] Susan J. Triggs, Ralph G. Powlesland, and Charles H. Daugherty. Genetic variation and conservation of kakapo (Strigops habroptilus: Psittaciformes). Conservation Biology, 3(1):92–96, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:1989:BCB

[TS89] Jack Ward Thomas and Hal Salwasser. Bringing conservation biology into a position of influence in natural resource management. Conservation Biology, 3(2):123–127, June 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorne:1988:DES

[TW88] E. Tom Thorne and Elizabeth S. Williams. Disease and endangered species: The black-footed Ferret as a recent example. Conservation Biology, 2(1):66–74, March 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vermeij:1989:IEL

[Ver89a] Geerat J. Vermeij. Invasion and extinction: The last three million years of North Sea pelecypod history. Conservation Biology, 3 (3):274–281, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vermeij:1989:SSW

[Ver89b] Geerat J. Vermeij. Saving the sea: What we know and what we need to know. Conservation Biology, 3(3):240–241, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vasquez:1989:UMF

[VG89] Rodolfo Vasquez and Alwyn H. Gentry. Use and misuse of forestharvested fruits in the Iquitos area. Conservation Biology, 3(4):350– 361, December 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vanVuren:1989:GCF

[vVH89] Dirk van Vuren and Philip W. Hedrick. Genetic conservation in feral populations of livestock. Conservation Biology, 3(3):312–317, September 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waller:1988:SRC

[Wal88] Donald M. Waller. Sharing responsibility for conserving diversity: The complementary roles of conservation biologists and public land agencies. Conservation Biology, 2(4):398–401, December 1988. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:1989:PBC

[WB89] Scott D. Wilson and Joyce W. Belcher. Plant and bird communities of native prairie and introduced Eurasian vegetation in Manitoba, Canada. Conservation Biology, 3(1):39–44, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitten:1987:OME

[WBNC87] Anthony J. Whitten, K. David Bishop, Stephen V. Nash, and Lynn Clayton. One or more extinctions born Sulawesi, Indonesia? Conservation Biology, 1(1):42–48, May 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitten:1987:ITP

[Whi87] Anthony J. Whitten. Indonesia’s transmigration program and its role in the loss of tropical rain forests. Conservation Biology, 1(3): 239–246, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcox:1987:E

[Wil87a] Bruce A. Wilcox. Editorial. Conservation Biology, 1(3):188–190, October 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:1987:LTR

[Wil87b] Edward O. Wilson. The little things that run the world (the importance and conservation of invertebrates). Conservation Biology, 1(4):344–346, December 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiles:1989:PEF

[WLP89] Gary J. Wiles, Thomas O. Lemke, and Nicholas H. Payne. Population estimates of fruit bats (Pteropus mariannus) in the Mariana Islands. Conservation Biology, 3(1):66–76, March 1989. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waller:1987:GVE

[WOG87] D. M. Waller, D. M. O’Malley, and S. C. Gawler. Genetic variation in the extreme endemic Pedicularis furbishiae, (Scrophulariaceae). Conservation Biology, 1(4):335–340, December 1987. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahner:1988:CWC

[Yah88] Richard H. Yahner. Changes in wildlife communities near edges. Conservation Biology, 2(4):333–339, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zwank:1988:FDB

[ZGD88] Phillip J. Zwank, James P. Geaghan, and Donna A. Dewhurst. Foraging differences between native and released Mississippi sandhill cranes: Implications for conservation. Conservation Biology, 2 (4):386–390, December 1988. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zwank:1987:SCP

[ZW87] Phillip J. Zwank and Charlotte D. Wilson. Survival of captive, parent-reared Mississippi sandhill cranes released on a refuge. Conservation Biology, 1(2):165–168, August 1987. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).